# Unified Multi-Modal Interleaved Document Representation for Information Retrieval

### v1

@@ 3.3 Reranker
Figure 2: Overview of IDentIfy. (a): In our document retriever, a query encoder represents a query (purple), and sections are encoded with a section encoder whose embeddings averaged to form a document representation (blue). Contrastive learning loss (red) is used for training the document retriever. (b): Reranker scores query-section relevance with the concatenation of the query and section, trained using Binary Cross-Entropy loss.
To enable fine-grained retrieval within documents beyond the retrieval of documents themselves, we introduce a section-level reranking mechanism that identifies the section most relevant to the input query. In particular, once the document is retrieved, the objective of the reranker $f_{\mathrm{R}}$ is to pinpoint the specific sections within the document that best match the query. We also note that this reranker is similarly operationalized with a single VLM along with a binary classifier on top of it, which directly measures the relevance of each query-section pair, illustrated in in Figure 2 (b).
@@
To train this reranker, we use the Binary Cross-Entropy (BCE) loss, formalized as follows:
$\begin{split}\mathcal{L}_{\text{reranker}}\!=\!\sum_{i=1}^{B}\sum_{j=1}^{S_{i}}\!\frac{1}{BS_{i}}\ell\left(\mathbf{y}_{(\mathrm{S}_{i,j})},\;f_{\mathrm{R}}\left(\left[{\bm{q}},\;{\bm{s}}_{i,j},\;\texttt{[EoS]}\right]\right)\right),\;\ell\left(y,\hat{y}\right)\!=\!-\left[y\log\hat{y}\!+\!(1\!-\!y)\log(1\!-\!\hat{y})\right],\end{split}$
where $S_{i}$ is the number of sections in the $i$-th document, $\mathbf{y}_{(\mathrm{S}_{i,j})}$ is the label for the $j$-th section of the $i$-th document ${\bm{s}}_{i,j}$ (with its value of one if relevant to the query ${\bm{q}}$ and zero otherwise), and $B$ is the batch size during training. Also, in this training process, the sections not labeled as relevant to the query are considered negative samples. Then, by minimizing $\mathcal{L}_{\text{reranker}}$, the reranker learns to predict section relevance for any query, thus refining our overall retrieval process by allowing the retrival of not just whole documents but also their most relevant sections, for multiple use cases of IR.
@@ 4 Experiments
4.1 Datasets
To evaluate the effectivenss of IDentIfy, we focus on multimodal IR tasks that require understanding of both the textual and visual cues within queries and documents, which align well with our goal of enhancing retrieval of multimodal interleaved documents. The datasets considered are as follows:
Encyclopedic-VQA (Mensink et al., 2023) is a large-scale visual question-answering (VQA) benchmark dataset, widely used for measuring the performance of multimodal IR models. Each query is linked to a specific section of a Wikipedia document (containing an answer for it) and is manually annotated by humans. Also, this dataset offers both text-only and multimodal queries. In addition to this, the queries are related to fine-grained properties of species and landmarks. Our experiments focus on the single-hop category where questions that can be answered in a single retrieval step.
ViQuAE (Lerner et al., 2022) is a dataset focused about human entities. It provides both text-based and multimodal queries, with each query linked to a specific section of a Wikipedia document that contains an answer (annotated by humans), which makes it an idea benchmark for section retrieval.
Open-WikiTable (Kweon et al., 2023) is an extension of WikiSQL (Zhong et al., 2017) and WikiTableQuestions (Pasupat & Liang, 2015), designed for open-domain table question answering that requires retrieval of the most relevant table from a broader corpus. For our experiments, we adapt this dataset, aiming at identifying the document or document section containing the target table, and correspondingly, utilize the WikiTableQuestions subset of Open-WikiTable that has labels for it.
4.2 Implementation Details
Model Training and Evaluation
We use LLaVA-NeXT-Interleave (Li et al., 2024) of 0.5B parameters as the basis VLM, for both the retriever and reranker. To take the advantage of larger batch sizes (while reducing GPU memory usage), we apply LoRA (Hu et al., 2022). Also, to further optimize the GPU usage, we combine four images into one, scaling each down to half of its original height and width. During retriever and reranker training, we consider four sections per document in representing documents and selecting negative samples. In contrast, during inference, we consider all sections within each document. For section retrieval, the top 25 documents retrieved are split into sections and passed to the rerankers. All experiments are conducted using a single H100 GPU.
Baselines
We compare our approach against a variety of IR baselines designed to capture different document representations. We start with Entity and Summary baselines, which are trained to retrieve documents based on their titles and summary sections. Next, we consider the Text-document retriever, which retrieves documents based on their textual content. Additionally, to consider a visual component, we consider the Single-image baseline, incorporating the first document image.
Evaluation Metrics
We evaluate the performance of the retriever and reranker with standard metrics: Recall@K (R@K) and Mean Reciprocal Rank@K (MRR@K). First, R@K measures whether the relevant document or section is retrieved within the top-K results. MRR@K evaluates the ranking quality by measuring the position of the first relevant item among the top-K retrieved results.
4.3 Results and Discussion
Interleaved format improves document retrieval.
Table 1: Comparison results of the document formats for document retrieval with multimodal queries. Entity uses the document title, Summary uses its first summary section, and Text-document uses only textual content. Based on the Text-document, we add a single image (+ Single-image) or interleaved multimodal content (+ Interleaved).
@@ 4.3 Results and Discussion
Table 2: Investigation of retrieval granularity in information retrieval. Two granularities are tested on section retrieval with multimodal queries: Passage splits a document by section boundaries, and, using the split passages as retrieval targets, trains the retriever; meanwhile, Document uses documents as retrieval targets. The same reranker is applied to items retrieved by each method. * indicates a result that does not leverage reranking.
@@
First, we report the retrieval performance on the Encyclopedic-VQA dataset in Table 2, where each query consists of both image and text content. From this, we observe that our approach achieves the best performance, with R@1 score improvements of 53.0%, 64.0%, and 25.0% compared to the Summary, Text-document, and Single-image retrieval baselines, respectively. The MRR@10 score similarly shows significant gains, with improvements of 36.1%, 48.5%, and 16.2% over the same baselines. This demonstrates the effectiveness of our approach in incorporating the interleaved multimodal format for document representations.
To further understand the source of these performance gains, we explore two levels of retrieval granularity: passages and documents. Specifically, the passage retriever uses individual sections of documents as retrieval units, while the document retriever treats entire documents as single units. Both models are trained on the Encyclopedic-VQA dataset for multimodal retrieval. Then, we use the same reranker to both sets of results from passage and document retrievers, to directly compare their performance. In Table 2, we observe that relying solely on the passage retriever (Passage*) results in suboptimal retrieval performance, highlighting the challenge in pinpointing the most relevant section within a document using traditional retrieval methods. In contrast, when the reranker is used alongside the document retriever, the performance significantly surpasses that of the passage retrieval, achieving a 22.7% improvement in R@1 and a 29.2% improvement in MRR@10, even though the document retriever provides eight times fewer retrieval units to the reranker. These results confirm the importance of leveraging holistic context from multiple, interrelated sections within documents. In addition to this, these findings also demonstrate the notable advantages of using the interleaved multimodal elements within documents, emphasizing the potential of this direction.
Interleaved format enhances document retrieval across modalities.
Table 3: Performance in document retrievals. (a): Results of retrieval for multimodal queries on InfoSeek and ViQuAE. (b): Results of retrieval for textual queries on Encyclopedic-VQA (Enc-VQA) and ViQuAE.
@@
Foramt Dataset R@1 R@10 R@100 MRR@10 Text-document InfoSeek 6.8 23.6 52.5 11.2 + Interleaved 10.2 30.4 57.3 15.7 Text-document ViQuAE 13.5 40.4 67.4 20.9 + Interleaved 17.5 46.0 69.4 26.3
@@
Table 4: Performance of section retrievals. (a): Results of retrieval for multimodal queries on Encyclopedic-VQA (Enc-VQA) and ViQuAE. (b): Results of retrieval for textual queries on Enc-VQA and ViQuAE. The same document retrieval outcomes are used in section retrieval to solely measure the reranker’s performance.
(a) Section Retrieval with Multimodal Queries
@@
(b) Section Retrieval with Textual Queries
@@
We further expand our experiments to two additional IR datasets, the InfoSeek and ViQuAE. As shown in Table 3 (a), our proposed retriever consistently surpasses the Text-document baseline in document retrieval with multimodal queries. Specifically, this leads to 50.0% and 29.6% improvements in the R@1 score, and 40.2% and 25.7% improvements in the MRR@10 score for the InfoSeek and ViQuAE, respectively. We also examine the impact of interleaved documents on textual retrieval tasks, where queries consist solely of text, and report the results in Table 3 (b). Then, the results demonstrate that the interleaved format offers advantages in retrieval of textual queries as well, resulting in 4.3% and 1.3% improvements in the R@1 score and 3.0% and 1.1% improvements in the MRR@10 score for the Encyclopedic-VQA and ViQuAE, respectively. We attribute these gains to the integration of multimodal content within documents, enabling the VLM to capture the multimodal alignment and to exploit its pre-existing knowledge for more effective document representations (Xu et al., 2024).
Interleaved format is also beneficial in section retrieval.
Similarly, we evaluate the efficacy of our approach in section retrieval across both multimodal and textual queries, using the Encyclopedic-VQA and ViQuAE datasets. First, in section retrieval with multimodal queries, our model attains 4.2% improvement in the R@1 score and 3.3% improvement in the MRR@10 score for the Encyclopedic VQA, as shown in Table 4 (a). Similarly, in section retrieval with textual queries, our model achieves 2.3% and 7.5% improvements in the R@1 score and 1.8% and 4.9% improvements in the MRR@10 score for the Encyclopedic and ViQuAE datasets, as shown in Table 4 (b). Overall, the design of our Interleaved rerankers exhibit superior or comparable performance to the Text-document rerankers. However, since the rerankers assess query relevance using a single section, they may lack access to broader contextual information from a document, which limits the potential performance gain compared to the retrievers. Nonetheless, the multimodal content interleaved within documents improves the reranker’s ability to evaluate the relevance of the query to individual sections.
Information retrieval of tabular contents in interleaved documents is challenging.
Table 5: Document and section retrieval results for tables, where Zero-shot denotes a model finetuned on Encyclopedic-VQA but not trained on the target dataset. Finetuned refers to additional training of the model on the target dataset. (a): Results for tabular document retrieval on the Open-WikiTable (OWT) dataset. (b): Textual section retrieval results on the ViQuAE dataset and tabular section retrieval results on the OWT dataset. (c): Reranker accuracy (Acc@1) of a classification task that identifies the section containing the query-associated table within a gold document containing multiple tables. Random indicates random selection in the task.
@@
We explore a retrieval task for tabular data, whose goal is to select the document or section containing the target table relevant to the input query. Specifically, we use the Open-WikiTable dataset to train the retriever and reranker, and then compare these trained models (Finetuned) with the models trained on the Encyclopedic-VQA dataset (Zero-shot). Then, as shown in Table 5 (a), despite Open-WikiTable consisting of only 3.2k training samples, the Finetuned retriever achieves strong retrieval performance. Meanwhile, the Zero-shot retriever demonstrates only about half of the R@1 score and the MRR@10 score of the Finetuned retriever, though it remains competitive in R@100.
In contrast, the performance trends for the rerankers exhibit notable differences. The discrepancies in R@10 and R@20 scores between the Zero-shot and Finetuned retrievers, denoted in color red in Table 5 (b), are much more pronounced in the Open-WikiTable (table retrieval) experiments than the ones in the ViQuAE (text retrieval) experiments. This highlights a substantial difference between textual and tabular modalities, despite both being represented using word tokens. This suggests that these two modalities may require different handling for retrieval, which we leave as future work.
Notably, the R@1 scores for tabular section retrieval are significantly lower than those for textual section retrieval (Table 5 (b)). To better understand the difficulty of identifying the query-relevant table, we use the Open-WikiTable dataset and design a classification task. In this task, both Zero-shot and Finetuned rerankers are provided with a golden document — a document containing the target of the input query — and should identify the section that contains the target table. Notably, this setup isolates the reranker’s ability to locate the target within a golden document. To ensure that the difficulty of the task is accurately assessed, we focus on documents containing multiple tables. Then, as shown in Table 5 (c), the Zero-shot reranker performs similarly to a random selection, failing to find the correct section. This accounts for its low R@10 and R@20 scores in the tabular retrieval task (Table 5 (b)). In aggregate, while the Finetuned reranker shows improved performance, it still misclassifies nearly half of the tables due to the high similarity between tables within the same document. When combined with tables from other documents, this further complicates the task of identifying the exact query-relevant table, as shown in Finetuned reranker’s low score in Table 5 (b).
4.4 Further Analysis and Ablation
@@
Figure 3: Trade-off between retrieval performance and training cost.
Table 6: Impact of negative sample selection in the reranker training. Top-K selects the top-k retrieved sections from the retriever as negatives. In-batch uses negatives from other sections in the same batch. In-document selects negatives from sections within the document containing the positive section.
Negative R@1 R@10 R@20 MRR@10
Top-K 38.1 53.7 55.3 44.4
In-batch 39.5 53.8 55.4 45.0
In-document (Ours) 42.4 53.6 55.7 46.3
In Table 2 and Table 2, we observe that using the comprehensive multimodal content and enriched contextual information significantly improves document retrieval performance. Accordingly, we anticipate further improvements as more sections are gathered to represent the document, during training. To validate this, we measure the document retrieval performance with varying the number of sections per document on the InfoSeek dataset for training. The results shown in Section 4.4 then indicate that incorporating more sections raises the MRR@10 score from 7.5 to 15.7. However, this performance boost comes with a clear trade-off; as the number of sections increases, the retriever must process additional end-of-section tokens, leading to higher GPU memory consumption. To balance resource limitations and performance gains, we select four sections per document for all experiments.
We explore another method to improve IR effectiveness by leveraging the entire document. Specifically, we investigate the use of sections from the same document as negatives for reranker training, namely In-document. We compare this approach with traditional methods, including Top-K, which selects the top-K retrieved sections as negatives, and In-batch, which uses the positive sections for other samples in the same batch as negatives. After training rerankers with each method, we evaluate section retrieval on the Encyclopedic-VQA dataset. The results shown in Section 4.4 demonstrate that our In-document approach achieves superior R@1 and MRR@10 scores. This suggests that the use of sections from the same document as negatives provides natural, cost-effective advantages thanks to their high similarity to the positive section. However, it does not consistently outperform the other methods on the R@10 score. We hypothesize that this inconsistency may arise from the strengths of each method: the In-document approach excels at distinguishing sections from the same document, while Top-K and In-batch methods better differentiate sections from different documents.
BCE loss applied to each section produces the best reranker performance.
Table 7: Performance of different reranker designs. Contrastive follows the same training pipeline used for retrievers but here it uses sections for retrieval. Document+BCE concatenates the input query with multiple sections gathered from the same document and uses BCE loss to train the re-ranker. Section+BCE concatenates the query with each section of the document, and the re-ranker is trained with the BCE loss.
@@
[…]
Then, in Table 7, we compare the section retrieval performance of different reranker designs on the Encyclopedic-VQA dataset. We find that the Contrastive reranker performs the worst, indicating that directly concatenating the query with the section at the input level provides more effective clues for query-section relevance assessment. Notably, this observation is consistent with conventional reranker approaches. Further, the Document+BCE reranker underperforms compared to the Section+BCE reranker, likely due to training constraints. Specifically, while the evaluation phase uses all sections within each document, the training phase is limited to a maximum of four sections per document, with an average of eight sections per document. Such a mismatch may degrade the model’s performance. Building on concurrent discoveries (Jiang et al., 2024; Lee et al., 2024), addressing these training constraints will potentially open up new reranker designs that can better handle long, interleaved documents using VLMs, ultimately improving section retrieval performance.
Rerankers require much larger datasets than retrievers.

### v2

@@ 3.3 Reranker
To enable fine-grained retrieval within documents beyond the retrieval of documents themselves, we introduce a section-level reranking mechanism that identifies the section most relevant to the query. In particular, once the document is retrieved, the objective of the reranker $f_{\mathrm{R}}$ is to pinpoint the specific sections within the document that best match the query. We also note that this reranker is similarly operationalized with the VLM along with a binary classifier on top of it, which directly measures the relevance of each query-section pair (Figure 2 (b)).
@@
To train this reranker, we use the binary cross-entropy loss, formulated as follow:
$\mathcal{L}_{\text{reranker}}$ $\!=\!\sum_{i=1}^{B}\sum_{j=1}^{S_{i}}\!\frac{1}{BS_{i}}\ell\left(\mathbf{y}_{{\bm{s}}_{i,j}},\;f_{\mathrm{R}}\left(\left[{\bm{q}},\;\hat{{\bm{s}}}_{i,j}\right]\right)\right),$
$\ell\left(y,\hat{y}\right)$ $\!=\!-\left[y\log\hat{y}\!+\!(1\!-\!y)\log(1\!-\!\hat{y})\right],$ (2)
[…]
@@ 4 Experiments
Table 1: Results with different document formats for retrieval.
@@
Table 2: Results with different section retrieval strategies. Document (Ours) performs document retrieval and section reranking, whereas Passage performs section retrieval and reranking. * denotes the model without reranking.
@@
Table 3: Performance on document retrievals. (a): Results of document retrieval for multimodal queries on InfoSeek and ViQuAE. (b): Results of document retrieval for textual queries on Encyclopedic-VQA (Enc-VQA) and ViQuAE.
@@
Format Dataset R@1 R@10 R@100 MRR@10 Text-document InfoSeek 6.8 23.6 52.5 11.2 + Interleaved 10.2 30.4 57.3 15.7 Text-document ViQuAE 13.5 40.4 67.4 20.9 + Interleaved 17.5 46.0 69.4 26.3
@@
4.1 Experimental Setups
Datasets
We evaluate IDentIfy on four benchmark datasets designed for multimodal IR tasks that require understanding of both the textual and visual cues within queries and documents, as follows: Encyclopedic-VQA (Mensink et al., 2023) is a large-scale benchmark for multimodal Visual Question Answering (VQA) with queries linked to specific Wikipedia sections and includes both text-only and multimodal queries; InfoSeek (Chen et al., 2023) is a knowledge-intensive VQA dataset with multimodal questions generated from Wikidata triples that include diverse entities such as landmarks, animals, and food; ViQuAE (Lerner et al., 2022) involves both text-based and multimodal queries about human entities, linked to annotated Wikipedia sections, making it ideal for evaluating section retrieval; Open-WikiTable (Kweon et al., 2023) extends WikiSQL (Zhong et al., 2017) and WikiTableQuestions (Pasupat and Liang, 2015), targeting open-domain table QA by identifying documents or sections containing relevant tables. We provide more details on datasets in Appendix A.
Baselines
We compare our approach against diverse baselines that capture different document representations. First, the Entity and Summary baselines retrieve documents based on their titles and summary sections, respectively, leveraging high-level textual cues. Also, the Text-document retriever baseline utilizes the full textual content of documents for retrieval. We further include the Single-image baseline that additionally leverages the first image of each document. IDentIfy is our model that holistically represents multimodal content (text, images, and tables) within documents.
Evaluation Metrics
To evaluate our approach, we use standard metrics: Recall@K (R@K) measures whether the relevant document or section appears within the top-K results; MRR@K measures how early the first relevant item is ranked (within top-K) by averaging its inverse rank across queries.
Implementation Details
We use LLaVA-NeXT-Interleave (Li et al., 2024b) of 0.5B parameters as the basis VLM for both the retriever and reranker. During training, documents are represented using randomly selected four sections, while in inference, we consider all sections within each document. For section-level retrieval, all sections within the top 25 retrieved documents are reranked. Experiments are conducted on a single H100 GPU.
4.2 Experimental Results and Analyses
Main Results
We report retrieval performance on the Encyclopedic-VQA dataset in Table 2, where queries include both text and images. We observe that IDentIfy achieves the best performance, improving R@1 scores by 53.0%, 64.0%, and 25.0% over Summary, Text-document, and Single-image retrieval baselines, respectively, with similar trends observed for other metrics. These results demonstrate the effectiveness of integrating multimodal content holistically into a unified representation. To further illustrate the advantages of our approach, we provide case studies in Appendix C.
We further examine the impact of our document retrieval and section reranking pipeline. In Table 2, the passage retriever represents individual sections as separate retrieval units, whereas the document retriever (ours) aggregates multiple section representations into a single representation. Then, we perform reranking over the retrieved sections or the sections from the retrieved documents, and then report the results in Table 2 (where * denotes the model without reranking). From this, we observe that the passage retriever without reranking (Passage*) achieves suboptimal retrieval performance, highlighting the challenge in pinpointing the most relevant section within a document using traditional retrieval methods. In contrast, when the reranker is used alongside the document retriever, the performance significantly surpasses that of the passage retrieval. These results confirm the importance of leveraging holistic context from multiple, interrelated sections within documents.
Interleaved format enhances document retrieval across modalities.
Table 4: Performance on section reranking. (a): Results of section reranking for multimodal queries on Encyclopedic-VQA (Enc-VQA) and ViQuAE. (b): Results of section reranking for textual queries on Enc-VQA and ViQuAE. For fair comparison between different methods, the reranking target (i.e., candidate sections from retrieved documents) are the same across them.
(a) Section Reranking with Multimodal Queries
@@ 4.2 Experimental Results and Analyses
(b) Section Reranking with Textual Queries
@@
Table 5: Retrieval results for tables, where Zero-shot denotes a model trained on Encyclopedic-VQA but not on the target dataset. Finetuned refers to additional training of the model on the target dataset. (a): Results for tabular document retrieval on Open-WikiTable (OWT). (b): Textual and tablular section retrieval results on ViQuAE and OWT datasets, respectively. (c): Reranker accuracy of a classification task that identifies the section containing the query-associated table given a gold document.
@@
We further expand our experiments to two additional datasets, InfoSeek and ViQuAE, and report document retrieval results. As shown in Table 3, our model consistently outperforms the Text-document baseline for both multimodal and text-only queries. We attribute these gains to the integration of multimodal content, allowing the VLM to capture richer alignments and leverage pre-existing knowledge for more effective document representation (Xu et al., 2024).
Interleaved format is also beneficial in section retrieval.
Similarly, we evaluate section retrieval performance on Encyclopedic-VQA and ViQuAE datasets, for both multimodal and textual queries. As shown in Table 4, our model outperforms the Text-document baseline in most cases. However, the performance gains over the baseline are smaller compared to the document retrieval setup. This is likely because section reranking focuses on evaluating the relationship between a single section and a query (rather than leveraging the holistic context of the entire document), and individual sections may lack the diverse multimodal information necessary for fully capturing the intent of queries.
Retrieving tables interleaved within documents is challenging.
We explore the retrieval task for tabular data, aiming to identify documents or sections containing query-relevant tables, and compare models trained on Encyclopedic-VQA (Zero-shot) with those additionally trained on Open-WikiTable (Finetuned). As shown in Table 5 (a), the Finetuned retriever outperforms the Zero-shot retriever on retrieving documents containing query-relevant tables. However, more fine-grained section reranking results (identifying sections containing query-relevant tables) in Table 5 (b) may reveal a notable modality-specific challenge: the performance of Zero-shot and Finetuned rerankers is considerably lower on table retrieval compared to their performance on text retrieval, despite both the text and tables being represented with word tokens. To better understand this, we design a classification task, where rerankers are tasked with identifying the correct section containing the target table within the golden document. Then, as shown in Table 5 (c), the Zero-shot reranker performs comparably to random selection, while the Finetuned reranker shows modest improvements. These findings highlight the intrinsic challenge of tabular retrieval, suggesting the need for table-specific modules to more holistically represent multimodal interleaved documents.
@@
Figure 3: Trade-off between performance (MRR@10) and training cost (GPU Memory) for retrieval.
Table 6: Comparison of negative sample selection strategies for reranker training: Top-K (top-k retrieved sections), In-batch (sections from other samples in the batch), and In-document (sections in the same document).
Negative R@1 R@20 MRR@10
Top-K 38.1 55.3 44.4
In-batch 39.5 55.4 45.0
In-document (Ours) 42.4 55.7 46.3
Table 7: Comparison of different training objectives for the reranker: Contrastive considers sections as retrieval units and uses the one for document retriever training; Document + BCE concatenates the query with multiple sections from the same document and uses the BCL loss; Section + BCE trains the reranker by concatenating the query with each section individually.
@@
To see how the number of sections used for representing each document impacts performance, we evaluate document retrieval on the InfoSeek dataset by varying the sections per document during training. As shown in Section 4.2, incorporating more sections improves MRR@10 from 7.5 to 15.7 due to leveraging richer multimodal and contextual information. However, this comes at the cost of increased computational requirements, as processing more sections raises GPU memory consumption.
In training the reranker, we investigate whether considering sections from the same document as negative examples (called In-document) is effective than other strategies, such as Top-K negatives (top-K retrieved sections based on their similarity with the input query) and In-batch negatives (positive sections from other samples in the same batch). As shown in Section 4.2, we observe that the In-document approach achieves superior performance especially on R@1, demonstrating its ability to effectively identify the most pertinent section among highly similar sections within the same document, i.e., its training objective can encourage the reranker to focus on fine-grained distinctions between closely related sections (within the same document).
BCE loss is the most effective to train the section reranker.
In our reranker design, we use a binary cross-entropy (BCE) loss by concatenating the query with each document section individually (Section + BCE), allowing the model to directly assess query-section relevance. As an alternative, we also explore a contrastive loss (Contrastive), which models section reranking similarly to document retrieval but uses sections as the retrieval units, and a variant of BCE loss (Document + BCE), where the query is concatenated with multiple sections (both positive and negative) from the same document. As shown in Table 7, the Section + BCE reranker outperforms both alternatives. Specifically, contrastive loss performs the worst, suggesting that direct concatenation of query and section provides clearer relevance signals, consistent with conventional reranking approaches. Moreover, while Document + BCE leverages inter-section context, its performance might be hindered by training constraints as the model processes fewer sections during training (Jiang et al., 2024a; Lee et al., 2024), and addressing it would be interesting future work.
Our Approach is Robust with Different Base Models.
Table 8: Results with another base model (LLaVA-OneVision) for document retrieval (with different document formats).
Format R@1 R@10 R@100 MRR@10
Entity 2.3 10.3 29.7 4.3
Summary 7.6 24.7 55.7 12.0
Text-document 7.0 24.1 50.4 11.7
+ Single-image 9.3 31.4 61.9 15.4
+ Interleaved (Ours) 12.1 36.1 62.5 18.2
To ensure the effectiveness of our approach across VLMs, we evalulate its performance with another VLM, LLaVA-OneVision (Li et al., 2024a), with 0.5 billion paramters, in addition to LLaVA-NeXT-Interleave (Li et al., 2024b) used in our main experiments. Results in Table 8 show that ours continues to outperform baselines, achieving a notable 30.1% gain in R@1 over the best baseline.
5 Conclusion
In this paper, we introduced IDentIfy, a novel IR framework designed to address the limitations of conventional methods that rely on textual content of documents and their segmented passages. Specifically, our approach sits on top of recent VLMs, which enables integration and representation of diverse multimodal content (including text, images, and tables) into a unified document representation. Also, unlike previous strategies that segment documents at the passage level, our method merges these segments to maintain the document’s structural coherence, while further introducing a reranking strategy for precise identification of relevant sections. Extensive experiments across various IR datasets demonstrated that IDentIfy consistently outperforms existing baselines, confirming that the interleaved multimodal representation significantly enhances the quality of the document retrieval. We believe IDentIfy represents a crucial step toward more comprehensive and contextually aware IR systems, capable of handling the increasing multimodality of modern information sources.
Limitations
Ethics Statement
In this work, we use a publicly available retrieval corpus for information retrieval tasks. However, the retrieval corpus may contain private, harmful, or biased content. Such undesirable features could unintentionally be reflected in the behavior of retrievers and rerankers trained on this data, potentially leading to ethical concerns during real-world deployment. However, current information retrieval techniques, including ours, do not address the retrieval of undesirable content. We recognize the critical need for safeguards to mitigate this issue. This is essential to ensure that information retrieval systems are reliable, fair, and safe for deployment.
Appendix A Details of Experimental Setups
Dataset configuration
Table 9 summarizes the key properties of the datasets used in our experiment, including query modality, target item, entity domain, number of entities, and whether a section ID is provided to indicate the section containing the answer. Additionally, we provide the number of samples in the training, evaluation, and test splits, as well as the size of the corpus. We provide a more detailed explanation of the datasets below.
  • Encyclopedic-VQA (Mensink et al., 2023) is a large-scale visual question-answering (VQA) benchmark dataset, widely used for measuring the performance of multimodal IR models. Each query is linked to a specific section of a Wikipedia document (containing an answer for it) and is manually annotated by humans. Also, this dataset offers both text-only and multimodal queries. In addition to this, the queries are related to fine-grained properties of species and landmarks. Our experiments focus on the single-hop category where questions can be answered in a single retrieval step.
  • ViQuAE (Lerner et al., 2022) is a dataset focused about human entities. It provides both textual and multimodal queries, with each query linked to a specific section of a Wikipedia document that contains an answer annotated by humans, which makes it an ideal benchmark for section retrieval.
  • Open-WikiTable (Kweon et al., 2023) is an extension of WikiSQL (Zhong et al., 2017) and WikiTableQuestions (Pasupat and Liang, 2015), designed for open-domain table question answering that requires retrieval of the most relevant table from a broader corpus. For our experiments, we adapt the WikiTableQuestions subset of Open-WikiTable, aiming at identifying the document or document section containing the target table.
Dataset pre-processing
In our study, we leverage interleaved multimodal content from Wikipedia documents. However, existing corpora associated with IR datasets often lack this content, typically only including the first few words of each document. Therefore, we download the HTML file of each Wikipedia document for corpus augmentation.
Since downloading contents for all documents across datasets is time- and memory-intensive, we preprocess subsets of each corpus, including documents relevant to queries in the training, evaluation, and test splits, along with unrelated documents.
Implementation Details
To take advantage of larger batch sizes (while reducing GPU memory usage), we apply LoRA (Hu et al., 2022). Also, to further optimize the GPU usage, we scale each image down to half of its original height and width and then combine four scaled-down images into a single composite image. All experiments are conducted using a single H100 GPU.
Appendix B Additional Experimental Results
Data Requirements for Models